NEED Question #2

1. U(x, y) = x1ax2(1-a)

a. Solve for the marshallian demands for x1 and x2, as functions of p1, p2, and m. (Hint: your solutions will be equations, not numbers).

b. For x1 find the own-price elasticity and income elasticity.

c. Suppose a = 0.2, m = 100, p1 = 2, and p2=8, find the quantities of x1 and x2.

d. What happens to these quantities when p1 doubles to $4?

e. What does this say about the price consumption curve (PCC)?

2. Suppose the price of one good increases. What is the substitution and the income effect of this price change? What else do you need to know to fully answer this?
Answer #1

1a a a a M aM Mshollion denand axt wce af am aft ladtiesty aM aM -1. Freeme eloobeity of X MB Incame Blarkterty= P a a-0.R m=Nhau P4 0 &X1OD 10 X2By maximizing utility subject to the budget constraint, we get the Marshallian Demand Functions. -a a 1-a MaximizeX ayoangen Can be written as POe ak axs 1-a ax -RAO a M-GM-B Soe s mtegftind a-1 a 7-a a -a a 1-a M=PX,+X2 a 1a a a a M aM Mshollion denand axt wce af am aft ladtiesty aM aM -1. Freeme eloobeity of X MB Incame Blarkterty= P a a-0.R m=